OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (anonymous):

slope is 4 so use the formula (y - y1) = m (x - x1)

OpenStudy (anonymous):

x – 4y = –18 4x – y = –18 x + 4y = 18 4x + y = 18

OpenStudy (anonymous):

4x+y=18